OpenStudy (anonymous):

(x)=-2x^2+2x+9 x-coordinate of the vertex y- coordinate symmetry min max of f(x) The value, f(1/2)=19/2 is Min or Max

hero (hero):

Not sure if I fully understand what's going on here 19/2 should be the max if you did this right

OpenStudy (anonymous):

You can find the x-coord of the vertex of a parabola using the formula \[x=-b/2a = -2/2(-2) = -1/2\]

OpenStudy (anonymous):

He may have a typo. why would the problem ask for f(+1/2) when the x-coordinate is x = /1/2?

hero (hero):

because x = 1/2, not -1/2

OpenStudy (anonymous):

Although it would still be a max (if the a = -2 is correct).

OpenStudy (anonymous):

I agree but that is what it says

OpenStudy (anonymous):

My fault! Sorry.

OpenStudy (anonymous):

Hero is correct (i made a sign error). The vertex is a max because the orientation of the parabola is opening downward. Evaluate the poly at x=1/2 to get the y-coordinate.

OpenStudy (anonymous):

y =f(1/2) = 18/2 = 9.5